# Optimal Non-Adaptive Group Testing with One-Sided Error Guarantees

### v3

@@ III Proofs
Here we provide the proofs of Theorem 2, Corollary 1, and Theorem 3, including a description of the algorithm mentioned in Theorem 2.
@@ III-A Proof of Theorem 2 (Achievability for SUBSET)
Since we are seeking to establish a rate of 1, we assume throughout the proof that the number of tests is $T=\big(k\log_{2}\frac{n}{k}\big)(1+\varepsilon)$ for arbitrarily small $\varepsilon>0$. We also implicitly condition on a fixed choice of $S$ (say $s:=\{1,\dotsc,k\}$, though we still adopt the original notation $S$ for consistency); the analysis will hold for any such choice, with the randomness only being over the test design (whose distribution is described below).
@@
where the equivalence holds due to both sets having the same size.
The tests follow a Bernoulli test design, where each item is included in each tests independently with some probability $p$. The value of $p$ is set to $p=\frac{\log 2}{k}\big(1+o(1)\big)$ such that each test is positive with probability $\frac{1}{2}$, thus maximizing the entropy of each test. The approximate recovery algorithm then checks through every $\bar{s}\in{\mathcal{S}}$, and looks for a unique choice such that, for all partitions $(s_{\mathrm{dif}},s_{\mathrm{eq}})$ of $\bar{s}$ satisfying $\lvert s_{\mathrm{dif}}\rvert>\beta k$, the following holds:
@@
where $\gamma_{\lfloor\beta k+1\rfloor},\gamma_{\lfloor\beta k+2\rfloor},\dots,\gamma_{k}$ are positive constants and $\imath^{T}\left({\mathsf{X}}_{s_{\mathrm{dif}}};\mathbf{y}\mid{\mathsf{X}}_{s_{\mathrm{eq}}}\right)$ is a conditional information density (the precise definition is omitted as it is not important for our purposes). The proof of [18, Theorem 3] reveals that for appropriate choices of $\gamma_{\lfloor\beta k+1\rfloor},\gamma_{\lfloor\beta k+2\rfloor},\dots,\gamma_{k}$ the algorithm succeeds in finding such a set if $\beta\in(0,1)$ is fixed and $T\geq\left(k\log_{2}\frac{n}{k}\right)(1+\varepsilon)$ for arbitrarily small $\varepsilon>0$. It turns out that their analysis can also be extended from constant $\beta$ to any $\beta=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ with $\lambda\in[0,\frac{1}{2})$, with only minor changes. The proof of this slightly strengthened statement is outlined in Appendix -B.
@@
A test $t\in\{1,2,\dots,T\}$ is explained by an item $i\in[n]$ if it holds that both ${\mathsf{X}}_{ti}=1$ and $y_{t}=1$, and there is no $t^{\prime}\in\{1,2,\dots,T\}$ such that ${\mathsf{X}}_{t^{\prime}i}=1$ and $y_{t^{\prime}}=0$.
@@
Input: Estimate $\hat{S}^{\prime}$ from the approximate recovery algorithm [18] with $\beta=\eta^{-}$, test matrix $\mathsf{X}$}, test outcome vector $\mathbf{y}$
@@
1: Initialize $T_{\mathrm{best}}=0,\ S_{\mathrm{best}}=\emptyset$
2:for all $\bar{S}$ such that $d_{\mathrm{H}}(\hat{S}^{\prime},\bar{S})\leq 3\eta^{-}k$ and $\lvert\bar{S}\rvert=(1-\eta^{-})k$ do
@@
The SUBSET algorithm first takes the estimate returned from the approximate recovery algorithm (which has size $k$), and then searches through all subsets of size $(1-\eta^{-})k$ within Hamming distance at most $3\eta^{-}k$ from the approximate recovery estimate, returning the set that explains the most tests. Note that this algorithm does not conduct any additional tests outside of the ones conducted to obtain $\hat{S}^{\prime}$. As such, the testing matrix $\mathsf{X}$} follows a Bernoulli test design with the following inclusion probability:
@@
The algorithm is successful if the returned set $\hat{S}$ is a subset of the defective set $S$. We will prove that this is indeed the case with high probability. To show this, it suffices to show that a subset of $S$ explains the most tests with high probability among all sets of size $(1-\eta^{-})k$ within distance $3\eta^{-}k$ of the approximate recovery estimate $\hat{S}^{\prime}$. To this end, a first thought would be to condition on the approximate recovery estimate $\hat{S}^{\prime}$ and analyze the sets around this estimate. However, this approach appears to be difficult, since conditioning on this event means that the entries of $\mathsf{X}$} are no longer independent, which complicates the analysis.
Instead, we study sets $S^{\prime}$ such that $d_{\mathrm{H}}(S,S^{\prime})\leq 5\eta^{-}k$, i.e., sets that lie within
${\mathcal{S}}_{\mathrm{an}}\coloneqq\{S^{\prime}\subseteq[n]:\lvert S^{\prime}\rvert=(1-\eta^{-})k,\ d_{\mathrm{H}}(S,S^{\prime})\leq 5\eta^{-}k\}.$ (12)
By comparison, the sets considered by the algorithm are as follows:
${\mathcal{S}}_{\mathrm{alg}}\coloneqq\{\bar{S}\subseteq[n]:\lvert\bar{S}\rvert=(1-\eta^{-})k,\ d_{\mathrm{H}}(\hat{S}^{\prime},\bar{S})\leq 3\eta^{-}k\}.$ (13)
We know that $d_{\mathrm{H}}(\hat{S}^{\prime},S)\leq 2\eta^{-}k$ (with high probability) due to the choice $\beta=\eta^{-}$ in Algorithm 1, and when this holds, the triangle inequality gives ${\mathcal{S}}_{\mathrm{alg}}\subseteq{\mathcal{S}}_{\mathrm{an}}$. In other words, if we prove a certain property for ${\mathcal{S}}_{\mathrm{an}}$, we can transfer that property to ${\mathcal{S}}_{\mathrm{alg}}$. When studying ${\mathcal{S}}_{\mathrm{an}}$, there is no need for conditioning of the kind discussed above.
Thus, we seek to show that some subset $\tilde{S}\subseteq S$ of size $(1-\eta^{-})k$ explains the most tests among those in ${\mathcal{S}}_{\mathrm{an}}$, while ensuring the choice of $\tilde{S}$ still is guaranteed to satisfy $\tilde{S}\in{\mathcal{S}}_{\mathrm{alg}}$. To do this, we consider making “local changes” so this suitably-chosen $\tilde{S}$, and show that with high probability, making these local changes reduces the number of tests explained. Before we do so, we introduce the following definition.
@@
This definition is relevant since when replacing an item $i$ with some other non-defective item, these “good” tests are no longer explained by defective items in the perturbed set. For a given $i\in S$, a test $t\in\{1,2,\dots,T\}$ is good with respect to $i$ with probability $\frac{\log 2(1+o(1))}{k}(1-\frac{\log 2(1+o(1))}{k})^{k-1}=\frac{\log 2}{2k}\big(1+o(1)\big)$ (see (11)). Hence, the number of tests that are good with respect to $i$ is distributed as $G_{i}\sim\mathrm{Bin}\big(T,\frac{\log 2}{2k}(1+o(1))\big)$. The following lemma gives a high probability lower bound on the number of items $i\in S$ such that $G_{i}$ is sufficiently large.
@@
If $T=\big(k\log_{2}\frac{n}{k}\big)(1+\varepsilon)$ for some $\varepsilon>0$, and $\eta^{-}=\Omega((\frac{k}{n})^{\lambda})$ for some $\lambda\in\big[0,\frac{1}{2}\big)$, [^2] We note that $k^{-\lambda(\frac{1}{\theta}-1)}=\Theta((\frac{k}{n})^{\lambda})$; the former parametrization is mainly used for the theorem statement while the latter is more convenient for this and subsequent proofs. then for any $\delta_{1}=1-o(1)$, it holds with probability 1o(1)1-o(1) that there are at least $k(1-\eta^{-})$ defective items $i$ such that $G_{i}\geq\frac{1-\delta_{1}}{2}\log\frac{n}{k}$.
@@
For a given $\smash{i\in S}$, the above calculations show that $\mu_{G}\coloneqq\mathbb{E}G_{i}=\frac{T\log 2}{2k}(1+o(1))\geq\frac{(1+\varepsilon)\log 2}{2}\log_{2}\frac{n}{k}(1+o(1))\newline \smash{=\frac{1+\varepsilon}{2}\log\frac{n}{k}\big(1+o(1)\big)}$. Hence, by standard binomial concentration bounds (Appendix -A) and the fact that $\mu_{G}\geq\frac{1}{2}\log\frac{n}{k}$, it holds that
$\mathbb{P}\left(G_{i}\leq\frac{1-\delta_{1}}{2}\log\frac{n}{k}\right)$ $\leq\mathbb{P}(G_{i}\leq(1-\delta_{1})\mu_{G})$ (14)
@@
where (18) follows since $(1-\delta_{1})\log(1-\delta_{1})+\delta_{1}=1-o(1)$ when $\delta_{1}=1-o(1)$. Let $Z=\sum_{i\in S}\mathds{1}\big\{G_{i}\leq\frac{1-\delta_{1}}{2}\log\frac{n}{k}\big\}$; by the linearity of expectation, $\mathbb{E}Z\leq k\left(\frac{k}{n}\right)^{\frac{1}{2}-o(1)}$, and by Markov’s inequality, it follows that
$\mathbb{P}\left(Z\geq\eta^{-}k\right)$ $\leq\frac{\mathbb{E}Z}{\eta^{-}k}$
$\leq\frac{\left(\frac{k}{n}\right)^{\frac{1}{2}-o(1)}}{\eta^{-}}$
@@
where $c$ is the value such that $\frac{1}{\eta^{-}}\leq c(\frac{n}{k})^{\lambda}$ for sufficiently large $n$ (since $\eta^{-}=\Omega((\frac{k}{n})^{\lambda})$), and (20) follows since $\lambda\in[0,\frac{1}{2})$ and $k=o(n)$. Hence, taking the contrapositive, we have with probability 1o(1)1-o(1) that at least $k-\eta^{-}k$ defective items $i\in S$ have $G_{i}\geq\frac{1-\delta_{1}}{2}\log\frac{n}{k}$ as claimed. ∎
@@
which follows from the definition of ${\mathcal{S}}_{\mathrm{alg}}$ in (13) along with $d_{\mathrm{H}}(\tilde{S},S)=|S\setminus\tilde{S}|=\eta^{-}k$, $d_{\mathrm{H}}(\hat{S}^{\prime},S)\leq 2\eta^{-}k$, and the triangle inequality. The importance of $\tilde{S}$ is highlighted in the following lemma.
@@
Conditioned on the event (21), and given $\tilde{S}$ specified in that event, the SUBSET algorithm succeeds provided that, for all $\ell=1,2,\dotsc,5\eta^{-}k$, and every possible $\tilde{S}^{\prime}$ formed by removing $\ell$ items from $\tilde{S}$ and replacing them with $\ell$ non-defective items, it holds that $\tilde{S}^{\prime}$ explains fewer tests than $\tilde{S}$.
Moreover, in order for such a set $\tilde{S}^{\prime}$ to explain at least as many tests as $\tilde{S}$, it must be the case that the $\ell$ (non-defective) items in $\tilde{S}^{\prime}\setminus\tilde{S}$ collectively explain at least $\frac{1-\delta_{1}}{2}\ell\log\frac{n}{k}$ tests.
@@
In view of the decoding rule in Algorithm 1, and the fact that $\tilde{S}\in\mathcal{S}_{\rm alg}$, in order to prove the first statement, it suffices to show that the sets $\tilde{S}^{\prime}$ formed in this manner (along with $\tilde{S}$ itself) collectively cover all of $\mathcal{S}_{\rm an}$ (and hence all of $\mathcal{S}_{\rm alg}$, since $\mathcal{S}_{\rm alg}\subseteq\mathcal{S}_{\rm an}$). The range $\ell=1,2,\dotsc,5\eta^{-}k$ suffices because the Hamming distance upon performing $\ell$ swaps is $2\ell$, and $\mathcal{S}_{\rm an}$ constrains the Hamming distance to be at most $10\eta^{-}k$ due to the triangle inequality (see (12)).
For the second claim, we note that by the final condition in (21), removing $\ell$ items from $\tilde{S}$ leads to at least $\frac{1-\delta_{1}}{2}\ell\log\frac{n}{k}$ tests no longer being explained. Thus, in order for $\tilde{S}^{\prime}$ to be favored over $\tilde{S}$ (i.e., explain more tests), it is necessary that the $\ell$ added non-defective items explain at least $\frac{1-\delta_{1}}{2}\ell\log\frac{n}{k}$ tests.
We proceed to analyze the probability of this being the case, starting with the following lemma.
@@
Let $\eta^{-}=\Theta((\frac{k}{n})^{\lambda})$ for some $\lambda\in(0,\frac{1}{2})$, and suppose that $\theta>\frac{1}{3}$. Then:
@@
    With probability 1o(1)1-o(1), under the choice $\delta_{2}=\sqrt{\frac{16\eta^{-}k\log\frac{1}{\eta^{-}}}{T}}$, it holds for all $\bar{s}\in{\mathcal{S}}_{\mathrm{an}}$ that the number of tests in which at least one item from $\bar{s}$ is present is at least $T(1-2^{-(1-\eta^{-})(1+o(1))})(1-\delta_{2})$.
@@
We now proceed with claim (ii). For a given test and $\bar{s}\in{\mathcal{S}}_{\mathrm{an}}$, at least one item $i\in\bar{s}$ is included in it with probability $1-\big(1-\frac{\log 2(1+o(1))}{k}\big)^{(1-\eta^{-})k}=1-e^{-(1-\eta^{-})\log 2(1+o(1))}=1-2^{-(1-\eta^{-})(1+o(1))}$. For brevity, we say that $\bar{s}$ is present in a test if at least one item from $\bar{s}$ is in the test. Thus, the number of tests in which $\bar{s}$ is present is distributed as $\mathrm{Bin}(T,1-2^{-(1-\eta^{-})(1+o(1))})$. Denoting the mean of this distribution as $\mu_{\bar{s}}$ and using the multiplicative Chernoff bound (Appendix -A), the number of tests in which $\bar{s}\in{\mathcal{S}}_{\mathrm{an}}$ is present is at most $\mu_{\bar{s}}(1-\delta_{2})$ with probability at most $\exp(-\frac{1}{2}\delta_{2}^{2}\mu_{\bar{s}})\leq\exp(-\frac{1}{8}\delta_{2}^{2}T)$, where we have used the crude bound $\mu_{\bar{s}}\geq\frac{1}{4}T$ for sufficiently large $n$ (with $\eta^{-}=o(1)$). Substituting the specified value of $\delta_{2}$, this bound becomes
$\mathbb{P}(\mathrm{Bin}(T,1-2^{-(1-\eta^{-})(1+o(1))})\leq(1-\delta_{2})\mu_{\bar{s}})\leq\exp\left(-2\eta^{-}k\log\frac{1}{\eta{-}}\right).$ (23)

### v4

@@ III Proofs
Here we provide the proofs of Theorem 2, Corollary 1, and Theorem 3, including a description of the algorithm mentioned in Theorem 2. We drop the superscripts from $\eta^{-},\eta^{+}$ throughout the proofs for ease of presentation; whether $\eta$ represents $\eta^{-}$ or $\eta^{+}$ will be clear from the context.
@@ III-A Proof of Theorem 2 (Achievability for SUBSET)
Preliminaries and Roadmap
Since we are seeking to establish a rate of 1, we assume throughout the proof that the number of tests is $T=\big(k\log_{2}\frac{n}{k}\big)(1+\varepsilon)$ for arbitrarily small $\varepsilon>0$. We also implicitly condition on a fixed choice of $S$ (say $s:=\{1,\dotsc,k\}$, though we often still adopt the original notation $S$); the analysis will hold for any such choice, with the randomness only being over the test design (whose distribution is described below).
@@
where the equivalence holds due to both sets having the same size. (Recall that the Hamming distance $d_{\mathrm{H}}$ between sets is interpreted as that between the associated length-$n$ indicator vectors.)
The tests follow a Bernoulli test design, where each item is included in each test independently with some probability $p$. The value of $p$ is set to $p=\frac{\log 2}{k}\big(1+o(1)\big)$ such that each test is positive with probability $\frac{1}{2}$, thus maximizing the entropy of each test. The approximate recovery algorithm then checks through every $\bar{s}\in{\mathcal{S}}$, and looks for a unique choice such that, for all partitions $(s_{\mathrm{dif}},s_{\mathrm{eq}})$ of $\bar{s}$ satisfying $\lvert s_{\mathrm{dif}}\rvert>\beta k$, the following holds:
@@
where $\gamma_{\lfloor\beta k+1\rfloor},\gamma_{\lfloor\beta k+2\rfloor},\dots,\gamma_{k}$ are positive constants and $\imath^{T}\left({\mathsf{X}}_{s_{\mathrm{dif}}};\mathbf{y}\mid{\mathsf{X}}_{s_{\mathrm{eq}}}\right)$ is a conditional information density (the precise definition is omitted as it is not important for our purposes). The proof of [28, Theorem 3] reveals that for appropriate choices of $\gamma_{\lfloor\beta k+1\rfloor},\gamma_{\lfloor\beta k+2\rfloor},\dots,\gamma_{k}$ the algorithm succeeds in finding such a set if $\beta\in(0,1)$ is fixed and $T\geq\left(k\log_{2}\frac{n}{k}\right)(1+\varepsilon)$ for arbitrarily small $\varepsilon>0$. It turns out that their analysis can also be extended from constant $\beta$ to any $\beta=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ with $\lambda\in[0,\frac{1}{2})$, with only minor changes. The proof of this slightly strengthened statement is outlined in Appendix -B.
@@
A test $t\in\{1,2,\dots,T\}$ is explained by an item $i\in[n]$ if it holds that both ${\mathsf{X}}_{ti}=1$ and $y_{t}=1$, and there is no $t^{\prime}\in\{1,2,\dots,T\}$ such that ${\mathsf{X}}_{t^{\prime}i}=1$ and $y_{t^{\prime}}=0$. Moreover, a test $t\in\{1,2,\dots,T\}$ is explained by a set ${\mathcal{J}}\subseteq[n]$ if $t$ is explained by at least one $i\in{\mathcal{J}}$.
@@
Input: Estimate $\hat{S}^{\prime}$ from the approximate recovery algorithm [28] with $\beta=\eta$, test matrix $\mathsf{X}$}, test outcome vector $\mathbf{y}$.
@@
1: Initialize $T_{\mathrm{best}}=0$
2:for all $\bar{S}$ such that $d_{\mathrm{H}}(\hat{S}^{\prime},\bar{S})\leq 3\eta k$ and $\lvert\bar{S}\rvert=(1-\eta)k$ do
@@
The SUBSET algorithm first takes the estimate returned from the approximate recovery algorithm (which has size $k$), and then searches through all subsets of size $(1-\eta)k$ within Hamming distance at most $3\eta k$ from the approximate recovery estimate, returning the set that explains the most tests. Note that this algorithm does not conduct any additional tests outside of the ones conducted to obtain $\hat{S}^{\prime}$. As such, the testing matrix $\mathsf{X}$} follows a Bernoulli test design with the following inclusion probability:
@@
The algorithm is successful if the returned set $\hat{S}$ is a subset of the defective set $S$. We will prove that this is indeed the case with high probability. To show this, it suffices to show that a subset of $S$ explains the most tests with high probability among all sets of size $(1-\eta)k$ within distance $3\eta k$ of the approximate recovery estimate $\hat{S}^{\prime}$. We will show this via the following high-level roadmap:
  1. 1.
[…]
  2. 2.
[…]
  3. 3.
    Finally, we show that the “goodness” property of $\tilde{S}$ outlined in Step 1 ensures that the event just described has sufficiently sufficiently small probability (even after a union bound over $\tilde{S}^{\prime}$ and $\ell$) via a careful argument based on binomial concentration.
Detailed Proof
We proceed to formalize the above outline. A first thought would be to condition on the approximate recovery estimate $\hat{S}^{\prime}$ and analyze the sets around this estimate. However, this approach appears to be difficult, since conditioning on this event means that the entries of $\mathsf{X}$} are no longer independent, which complicates the analysis.
Instead, we study sets $S^{\prime}$ such that $d_{\mathrm{H}}(S,S^{\prime})\leq 5\eta k$, i.e., sets that lie within
${\mathcal{S}}_{\mathrm{an}}\coloneqq\{S^{\prime}\subseteq[n]:\lvert S^{\prime}\rvert=(1-\eta)k,\ d_{\mathrm{H}}(S,S^{\prime})\leq 5\eta k\}.$ (12)
(The subscript “an” indicates being used for mathematical “analysis”.) By comparison, the sets considered by the algorithm are as follows:
${\mathcal{S}}_{\mathrm{alg}}\coloneqq\{\bar{S}\subseteq[n]:\lvert\bar{S}\rvert=(1-\eta)k,\ d_{\mathrm{H}}(\hat{S}^{\prime},\bar{S})\leq 3\eta k\}.$ (13)
We know that $d_{\mathrm{H}}(\hat{S}^{\prime},S)\leq 2\eta k$ (with high probability) due to the choice $\beta=\eta$ in Algorithm 1, and when this holds, the triangle inequality gives ${\mathcal{S}}_{\mathrm{alg}}\subseteq{\mathcal{S}}_{\mathrm{an}}$. In other words, if we prove a certain property for ${\mathcal{S}}_{\mathrm{an}}$, we can transfer that property to ${\mathcal{S}}_{\mathrm{alg}}$. When studying ${\mathcal{S}}_{\mathrm{an}}$, there is no need for conditioning of the kind discussed above.
Thus, we seek to show that some subset $\tilde{S}\subseteq S$ of size $(1-\eta)k$ explains the most tests among those in ${\mathcal{S}}_{\mathrm{an}}$, while ensuring that the choice of $\tilde{S}$ still satisfies $\tilde{S}\in{\mathcal{S}}_{\mathrm{alg}}$. To do this, we consider making “local changes” to this suitably-chosen $\tilde{S}$, and show that with high probability, making these local changes reduces the number of tests explained. Before we do so, we introduce the following definition.
@@
This definition is relevant since when replacing an item $i\in\tilde{S}$ by some other non-defective item, these “good” tests are no longer explained by defective items in the perturbed set. For a given $i\in S$, a test $t\in\{1,2,\dots,T\}$ is good with respect to $i$ with probability $\frac{\log 2(1+o(1))}{k}(1-\frac{\log 2(1+o(1))}{k})^{k-1}=\frac{\log 2}{2k}\big(1+o(1)\big)$ (see (11)). Hence, the number of tests that are good with respect to $i$ is distributed as $G_{i}\sim\mathrm{Bin}\big(T,\frac{\log 2}{2k}(1+o(1))\big)$. The following lemma gives a high probability lower bound on the number of items $i\in S$ such that $G_{i}$ is sufficiently large.
@@
If $T=\big(k\log_{2}\frac{n}{k}\big)(1+\varepsilon)$ for some $\varepsilon>0$, and $\eta=\Omega((\frac{k}{n})^{\lambda})$ for some $\lambda\in\big[0,\frac{1}{2}\big)$, [^2] We note that $k^{-\lambda(\frac{1}{\theta}-1)}=\Theta((\frac{k}{n})^{\lambda})$; the former parametrization is mainly used for the theorem statement while the latter is more convenient here and in subsequent proofs. then for any $\delta_{1}=1-o(1)$, it holds with probability 1o(1)1-o(1) that there are at least $k(1-\eta)$ defective items $i$ such that $G_{i}\geq\frac{1-\delta_{1}}{2}\log\frac{n}{k}$.
@@
For a given $\smash{i\in S}$, the above calculations show that
$\mu_{G}$ $\coloneqq\mathbb{E}G_{i}$ (14)
$=\frac{T\log 2}{2k}(1+o(1))$ (15)
$=\frac{(1+\varepsilon)\log 2}{2}\log_{2}\frac{n}{k}(1+o(1))$ (16)
$=\frac{1+\varepsilon}{2}\log\frac{n}{k}\big(1+o(1)\big).$ (17)
Hence, by standard binomial concentration bounds (Appendix -A) and the fact that $\mu_{G}\geq\frac{1}{2}\log\frac{n}{k}$, it holds that
$\mathbb{P}\Big(G_{i}$ $\leq\frac{1-\delta_{1}}{2}\log\frac{n}{k}\Big)\leq\mathbb{P}(G_{i}\leq(1-\delta_{1})\mu_{G})$ (18)
@@
where (22) follows since $(1-\delta_{1})\log(1-\delta_{1})+\delta_{1}=1-o(1)$ when $\delta_{1}=1-o(1)$. Let
$Z=\sum_{i\in S}\mathds{1}\left\{G_{i}\leq\frac{1-\delta_{1}}{2}\log\frac{n}{k}\right\};$ (23)
by the linearity of expectation, $\mathbb{E}Z\leq k\left(\frac{k}{n}\right)^{\frac{1}{2}-o(1)}$, and by Markov’s inequality, it follows that
$\mathbb{P}\left(Z\geq\eta k\right)$ $\leq\frac{\mathbb{E}Z}{\eta k}$
$\leq\frac{\left(\frac{k}{n}\right)^{\frac{1}{2}-o(1)}}{\eta}$
@@
where $c>0$ is a constant such that $\frac{1}{\eta}\leq c(\frac{n}{k})^{\lambda}$ for sufficiently large $n$ (since $\eta=\Omega((\frac{k}{n})^{\lambda})$), and (25) follows since $\lambda\in[0,\frac{1}{2})$ and $k=o(n)$. Hence, we have with probability 1o(1)1-o(1) that at least $k-\eta k$ defective items $i\in S$ have $G_{i}\geq\frac{1-\delta_{1}}{2}\log\frac{n}{k}$ as claimed. ∎
@@
$\exists$ $\tilde{S}\,:\,\big(\lvert\tilde{S}\rvert=(1-\eta)k\big)\wedge\big(\tilde{S}\subseteq S\big)$
@@
which follows from the definition of ${\mathcal{S}}_{\mathrm{alg}}$ in (13) along with $d_{\mathrm{H}}(\tilde{S},S)=|S\setminus\tilde{S}|=\eta k$, $d_{\mathrm{H}}(\hat{S}^{\prime},S)\leq 2\eta k$, and the triangle inequality. The importance of $\tilde{S}$ is highlighted in the following lemma.
@@
Conditioned on the event (26), and given $\tilde{S}$ specified in that event, the SUBSET algorithm succeeds provided that, for all $\ell=1,2,\dotsc,5\eta k$, and every possible $\tilde{S}^{\prime}$ formed by removing $\ell$ items from $\tilde{S}$ and replacing them with $\ell$ non-defective items, it holds that $\tilde{S}^{\prime}$ explains fewer tests than $\tilde{S}$.
Moreover, when any set of $\ell$ items ${\mathcal{D}}_{\ell}\subset\tilde{S}$ is removed from $\tilde{S}$, these removals collectively lead to at least $\frac{1-\delta_{1}}{2}\ell\log\frac{n}{k}$ tests no longer being explained by $\tilde{S}\setminus{\mathcal{D}}_{\ell}$.
@@
In view of the decoding rule in Algorithm 1, and the fact that $\tilde{S}\in\mathcal{S}_{\rm alg}$, in order to prove the first statement, it suffices to show that the sets $\tilde{S}^{\prime}$ formed in this manner (along with $\tilde{S}$ itself) collectively cover all of $\mathcal{S}_{\rm an}$ (and hence all of $\mathcal{S}_{\rm alg}$, since $\mathcal{S}_{\rm alg}\subseteq\mathcal{S}_{\rm an}$). The range $\ell=1,2,\dotsc,5\eta k$ suffices because the Hamming distance upon performing $\ell$ swaps is $2\ell$, and any two sets in $\mathcal{S}_{\rm an}$ have Hamming distance at most $10\eta k$ due to the triangle inequality (see (12)).
For the second claim, we note that by the final condition in (26), removing any item from $\tilde{S}$ leads to at least $\frac{1-\delta_{1}}{2}\log\frac{n}{k}$ tests no longer being explained. Thus, removing any set of $\ell$ items from $\tilde{S}$ leads to at least $\frac{1-\delta_{1}}{2}\ell\log\frac{n}{k}$ tests no longer being explained. ∎
We proceed to analyze the probability of some $\tilde{S}^{\prime}$ from Lemma 5 explaining more tests than $\tilde{S}$, starting with the following lemma.
@@
Fix $\lambda\in(0,\frac{1}{2})$, let $\eta=\Theta((\frac{k}{n})^{\lambda})$, and suppose that $\theta>\frac{1}{3}$. Then:
@@
    With probability 1o(1)1-o(1), under the choice $\delta_{2}=\sqrt{\frac{16(3+\frac{1}{\lambda})\eta k\log\frac{1}{\eta}}{T}}$, it holds for all $\bar{s}\in{\mathcal{S}}_{\mathrm{an}}$ that the number of tests in which at least one item from $\bar{s}$ is present is at least $T(1-2^{-(1-\eta)(1+o(1))})(1-\delta_{2})$.
@@
We now proceed with claim (ii). For a given test and $\bar{s}\in{\mathcal{S}}_{\mathrm{an}}$, at least one item $i\in\bar{s}$ is included in it with probability $1-\big(1-\frac{\log 2(1+o(1))}{k}\big)^{(1-\eta)k}=1-e^{-(1-\eta)\log 2(1+o(1))}=1-2^{-(1-\eta)(1+o(1))}$. For brevity, we say that $\bar{s}$ is present in a test if at least one item from $\bar{s}$ is in the test. Thus, the number of tests in which $\bar{s}$ is present is distributed as $\mathrm{Bin}(T,1-2^{-(1-\eta)(1+o(1))})$. Denoting the mean of this distribution as $\mu_{\bar{s}}$ and using the multiplicative Chernoff bound (Appendix -A), the number of tests in which $\bar{s}\in{\mathcal{S}}_{\mathrm{an}}$ is present is at most $\mu_{\bar{s}}(1-\delta_{2})$ with probability at most $\exp(-\frac{1}{2}\delta_{2}^{2}\mu_{\bar{s}})\leq\exp(-\frac{1}{8}\delta_{2}^{2}T)$, where we have used the crude bound $\mu_{\bar{s}}\geq\frac{1}{4}T$ for sufficiently large $n$ (with $\eta=o(1)$). Substituting the specified value of $\delta_{2}$, this bound becomes
$\mathbb{P}(\mathrm{Bin}(T,1$- $2^{-(1-\eta)(1+o(1))})\leq(1-\delta_{2})\mu_{\bar{s}})$
$\leq\exp\left(-2\Big(3+\frac{1}{\lambda}\Big)\eta k\log\frac{1}{\eta}\right).$ (28)